Calculate (a) the de Broglie wavelength of an electron moving with a velocity of 5.0x 105 ms–1 and (b) relative de Broglie wavelength of an atom of hydrogen and atom of oxygen moving with the same velocity (h = 6.63 x 10–34 kg m2 s–1)

Grade:11

1 Answers

Sakshi
8 years ago
Hello students..

As we know the de-Broglie wavelength can be written as
λ=h/mv
where h is the Planks constant, m is the mass of the electron and v is the velocity of electron.
On substituting the above values, one can find the value of de broglie wavelength.
For the part 2,
we can write,
λHO = (h*mO*vO)/(mH*vH*h) = (mO/mH) = 16/1=16:1

Hope you find the explaination worth!!!